OpenStudy (anonymous):

PROBLEM SOLVING USING SYSTEMS OF EQUATIONS Suppose you have $20 in your bank account. You start saving $5 each week. your friend has $5 in his bank account and saves $10 each week. Assume neither you or your friend make any withdraws. How many weeks will you and your friend have the same amount of money in your account, and how much money will each of you have?

OpenStudy (anonymous):

Let us assume that after n weeks you and your friend is going to have same amount. After n weeks the money in your account will be 20 (initial amt) + 5 (savings each week) * n or simply 20 + 5n Similarly, after n week your friend is going to have 5 + 10n As per the questions these two amounts are same, hence 20 + 5n = 5 + 10n => 5n = 15 => n = 3 (Desired number of weeks) hence, money in each account is 20 + 5*3 = $35 (Second part)

OpenStudy (anonymous):

im sorry i dont understand your explaination but thankyou so much i will give you a medal

OpenStudy (anonymous):

tell me which part dint you understand

OpenStudy (anonymous):

im a slow learner, so the two equations for starters. whatb are the two equations i will be working with?